OpenStudy (anonymous):

How many grams of carbon dioxide (CO2) are formed from 0.205 g of octane?

OpenStudy (abb0t):

This is assuming combustion, meaning: \(\sf \color{red}{C_8H_{18}+O_2 \rightarrow CO_2 + H_2O}\) NOW, balance it.

OpenStudy (anonymous):

2C8H18+25O2-->16CO2+18H2O Now what would I do?

OpenStudy (abb0t):

use diimensional analysis to convert from grams of octane to grams of CO\(_2\). Start: \(\sf 0.205~g~octance \times \frac{1 mol~octane}{mass~of~Octane} \times \frac{mol~CO_2}{mol~octane} \times \frac{mass~of~1~CO_2~molecule}{mol~of~CO_2} =\color{red}{mass~of~CO_2}\)

OpenStudy (abb0t):

\(\sf \cancel{0.205~g~octance} \times \frac{\cancel{1 mol~octane}}{\cancel{mass~of~Octane}} \times \frac{\cancel{mol~CO_2}}{\cancel{mol~octane}} \times \frac{mass~of~1~CO_2~molecule}{\cancel{mol~of~CO_2}} =\color{red}{mass~of~CO_2}\) notice how everything cancels out nicely.

OpenStudy (anonymous):

would it be .088g CO2 ?

OpenStudy (abb0t):

I do not have a calculator with me, but if you BALANCED your chemical equation and then followed what I showed you up there, then you should get the correct answer.

OpenStudy (abb0t):

I highly doubt that the ratio is 1:1 for octane to carbon dioxide.

OpenStudy (anonymous):

Okay thank you would you mind helping me set up another problem? AB reacts with C2 to yield AC and B. 2AB+C2 -->2AC+2B The molar mass of AB is 75.0 g/mol, the molar mass of AC is 60.0 g/mol, and the molar mass of B is 50.0 g/mol. What is the theoretical yield (in grams) of AC from the reaction of 29.0 g of AB? Thanks appreciate it!